Determine the maximum uniformly distributed load, a simply supported 40cm deep and 20cm wide timber beam can carry over a span of 4m. The maximum permissible bending stress for the timber is 2N/mm
1
Expert's answer
2020-08-14T16:01:39-0400

"I_x=\frac{bh^3}{12}=\frac{0.2\cdot0.4^3}{12}=0.00107(m^4)"


"\sigma_{max}=\frac{M_{max}c}{I_x}\to M_{max}=\sigma_{max}I_x/c="


"=2\cdot10^6\cdot0.00107/0.2=10700(N\cdot m)"


"M=2qx-qx^2\/2"


For "x=2(m)" "M=M_{max}"


"2q\cdot2-q\cdot2^2/2=10700\to q=5350(N/m)". Answer